OpenStudy (anonymous):

Jason considered two similar televisions at a local electronics store. The generic version was based off the brand name and was the size of the brand name. If the generic television set is 12 inches by 24 inches, what are the dimensions of the brand name television? 16 inches by 32 inches 9 inches by 18 inches 18 inches by 36 inches 8 inches by 16 inches

OpenStudy (valpey):

Same size meaning same area? WTF kind of TV is this?

OpenStudy (anonymous):

Jason considered two similar televisions at a local electronics store. The generic version was based off the brand name and was 2/3 the size of the brand name. If the generic television set is 12 inches by 24 inches, what are the dimensions of the brand name television? 16 inches by 32 inches 9 inches by 18 inches 18 inches by 36 inches 8 inches by 16 inches

OpenStudy (valpey):

Much better. 2/3 the area or each of the dimensions? 12 is 2/3 of 18 and 24 is 2/3 of 36.

OpenStudy (valpey):

The Area of the generic is 12*18 = 216 sq.in. which is 2/3 of 324 sq.in. but I don't think this is what they are asking.

OpenStudy (anonymous):

im not sure either this question is kind of confusing to me

OpenStudy (anonymous):

@ganeshie8 can you help lol?

ganeshie8 (ganeshie8):

Okay. here is the clue : Size of TV is measured across diagonal.

OpenStudy (anonymous):

Is the answer D?

ganeshie8 (ganeshie8):

\(\textbf{generic = } \frac{2}{3} \textbf{brand}\) \(=>\textbf{brand = } \frac{3}{2} \textbf{generic}\) so simply multiply generic dimensions with 3/2 to get brand dimensions 12 * 3/2 = 18 24 * 3/2 = ?

OpenStudy (anonymous):

36

ganeshie8 (ganeshie8):

yep :)

OpenStudy (anonymous):

Thankyouuu :D